Please help me with my last problem::
On cloudy days, the sun intensity I (in calories per square centimeter) is given by I = I(max) sin‚‚^2 (pi t / D) where I(max) is the largest sun intensity and D is the number of daylight hours. If D = 12, how many hours after sunrise is intensity exactly half of the maximum level? (Assume t is in between zero and D)
Thank you very much!!

Answer by stanbon(75887)

On cloudy days, the sun intensity I (in calories per square centimeter) is given by I = I(max) sin^2 (pi t / D) where I(max) is the largest sun intensity and D is the number of daylight hours. If D = 12, how many hours after sunrise is intensity exactly half of the maximum level? (Assume t is in between zero and D)
-----
Solve sin^2[(pi/12)t] = 1/2
---
sin[(pi/12)t] = sqrt(2)/2
---
(pi/12)t = pi/4
---
(t/12) = 1/4
---
t = 3 hrs.
====================
Cheers,
Stan H.